OpenStudy (anonymous):

Find (dy/dt): y=4(x^3)

OpenStudy (anonymous):

do you mean dy/dx?

OpenStudy (anonymous):

No, the question asks for (dy/dt).

OpenStudy (anonymous):

so, need more information

OpenStudy (anonymous):

agreed.

OpenStudy (anonymous):

That's all I was given.

sam (.sam.):

As a function of time that'd be zero

OpenStudy (asnaseer):

\[\frac{dy}{dt}=\frac{dy}{dx}\times\frac{dx}{dt}\]If \(\frac{dx}{dt}\) is zero, then \(\frac{dy}{dt}\) will also be zero

OpenStudy (asnaseer):

@madster007 - does that make sense to you?

OpenStudy (anonymous):

Yes! Thanks! :)

OpenStudy (asnaseer):

Good - I'm glad I was able to help :)